OpenStudy (anonymous):

Solve the equation (1/5)^(2-x) = 25? I know you need to get a (1/5) on the right side but how do you do that?

OpenStudy (angela210793):

(1/5)^(2-x) = 25 (1/5)^(2-x)=(1/5)^(-2) 2-x=-2 -x=-4 x=4 :)

OpenStudy (anonymous):

thank you!

OpenStudy (angela210793):

uw :)